40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Aug 2 11:37:38 2019, gautam, Update, ALS, EY IR ALS Assay IMG_7780.JPGEy_MM_20190802.pdf
    Reply  Tue Aug 6 23:09:20 2019, gautam, Update, ALS, EY table work 
       Reply  Fri Aug 9 16:37:39 2019, gautam, Update, ALS, More EY table work IMG_7791.JPGfiberCleaning.pdf
Message ID: 14838     Entry time: Fri Aug 9 16:37:39 2019     In reply to: 14835
Author: gautam 
Type: Update 
Category: ALS 
Subject: More EY table work 

Summary:

  1. 220 uW / 600 uW (~36 % mode-matching) of IR light coupled into fiber at EY.
  2. Re-connected the RF chain from the beat mouth output on the PSL table to the DFD setup at 1Y2.
  3. A beat note was found between the PSL and EY beams using the BeatMouth.

Motivation:

We want to know that we can lock the interferometer with the ALS beat note being generated by beating IR pickoffs (rather than the vertex green transmission). The hope is also to make the ALS system good enough that we can transition the CARM offset directly to 0 after the DRMI is locked with arms held off resonance.

Details:

Attachment #1: Shows the layout. The realized MM is ~36 %. c.f. the 85% predicted by a la mode. It is difficult to optimize much more given the tight layout, and the fact that these fast lenses require the beam to be well centered on them. They are reasonably well aligned, but I don't want to futz around with the pointing into the doubling crystal. Consequently, I don't have much control over the pointing.

Attachment #2: Shows pictures of the fiber tips at both ends before/after cleaning. The tips are now much cleaner.

The BeatMouth NF1611 DC monitor reports ~580 mV with only the EY light incident on it. This corresponds to ~60 uW of light making it to the photodiode, which is only 25% of what we send in. This is commensurate with the BS loss + mating sleeve losses.

To find the beat between PSL and EY beams, I had to change the temperature control MEDM slider for the EY laser to -8355 cts (it was 225 cts). Need to check where this lies in the mode-hop scan by actually looking at the X-tal temperature on the front panel of the EY NPRO controller - we want to be at ~39.3 C on the EY X-tal, given the PSL X-tal temp of ~30.61 C. Just checked it, front panel reports 39.2C, so I think we're good.

Next steps:

  • Fix the IMC suspension
  • Measure the ALS noise for the Y arm
  • Determine if improvements need to be made to the IR beat setup (e.g. more power? better MM? etc etc).

EY enclosure was closed up and ETMY Oplev was re-enabled after my work. Some cleanup/stray beam dumping remains to be done, I will enlist Chub's help on Monday.

Attachment 1: IMG_7791.JPG  1.690 MB  Uploaded Fri Aug 9 17:50:05 2019  | Hide | Hide all
IMG_7791.JPG
Attachment 2: fiberCleaning.pdf  72.892 MB  Uploaded Fri Aug 9 17:50:14 2019  | Hide | Hide all
fiberCleaning.pdf
ELOG V3.1.3-